# A universal black-box quantum Monte Carlo approach to quantum phase transitions

### v2

@@ E.2 Estimating $\langle H_{1}\rangle$ and $\langle H_{1}(\tau)H_{1}\rangle$
$\frac{\Lambda_{0}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=0}^{q}\Lambda_{0}(z_{j})e^{-\tau[E_{z_{j}},\ldots,E_{z_{q}}]}e^{-(\beta-\tau)[E_{z_{0}},\ldots,E_{z_{j}}]}$ =^$\mathrel{\widehat{=}}\langle\Lambda_{0}(\tau)\Lambda_{0}\rangle$ (97)
$\frac{\delta_{P_{k}}^{(q)}\Lambda_{k}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=0}^{q}\Lambda_{0}(z_{j})e^{-\tau[E_{z_{j}},\ldots,E_{z_{q-1}}]}e^{-(\beta-\tau)[E_{z_{0}},\ldots,E_{z_{j}}]}$ =^$\mathrel{\widehat{=}}\langle(\Lambda_{k}D_{k}P_{k})(\tau)\Lambda_{0}\rangle$ (98)
$\frac{\Lambda_{0}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=1}^{q-1}\delta^{(j)}_{P_{l}}\Lambda_{l}(z_{j})e^{-\tau[E_{z_{j}},\ldots,E_{z_{q}}]}e^{-(\beta-\tau)[E_{z_{0}},\ldots,E_{z_{j-1}}]}$ =^$\mathrel{\widehat{=}}\langle\Lambda_{0}(\tau)(\Lambda_{l}D_{l}P_{l})\rangle$ (99)
$\frac{\delta_{P_{k}}^{(q)}\Lambda_{k}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=1}^{q-1}{\delta_{P_{l}}^{(j)}}{\Lambda_{l}(z_{j})}e^{-\tau[E_{z_{j}},\ldots,{E_{z_{q-1}}}]}e^{-(\beta-\tau)[E_{z_{0}},\ldots,{E_{z_{j-1}}}]}$ =^$\mathrel{\widehat{=}}\langle(\Lambda_{k}D_{k}P_{k})(\tau)(\Lambda_{l}D_{l}P_{l})\rangle,$ (100)
where each estimator follows as different special cases of the derivation of Eq. 119 of Ref. [26]. As with the $\langle D_{0}(\tau)D_{0}\rangle$ estimator, each of these estimators requires $O(q^{2})$ operations, so the total complexity to evaluate $\langle H_{1}(\tau)H_{1}\rangle$ is simply $O(R^{2}q^{2})$.
Although these estimators are more complicated than the pure diagonal counterparts Section D.2, the big picture for ES and FS is the same. Namely, the only non-trivial $\tau$ dependence is contained in the divided differences of the correlator estimator, for which we can apply the divided difference integral relations we have already derived.
E.3 General ES estimator
@@
$\frac{\Lambda_{0}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=0}^{q}\Lambda_{0}(z_{j})e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}},E_{z_{j}}]}$ =^$\mathrel{\widehat{=}}\int_{0}^{\beta}\langle\Lambda_{0}(\tau)\Lambda_{0}\rangle\mathrm{d}\tau$ (101)
$\frac{\delta_{P_{k}}^{(q)}\Lambda_{k}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=1}^{q-1}{\delta_{P_{l}}^{(j)}}{\Lambda_{l}(z_{j})}e^{-\beta[E_{z_{0}},\ldots,E_{z_{q-1}}]}$ =^$\mathrel{\widehat{=}}\int_{0}^{\beta}\langle(\Lambda_{k}D_{k}P_{k})(\tau)(\Lambda_{l}D_{l}P_{l})\rangle\mathrm{d}\tau.$ (102)
@@ E.4 General FS estimator
$\frac{\Lambda_{0}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=0}^{q}\Lambda_{0}(z_{j})\sum_{r=0}^{j}e^{-\frac{\beta}{2}[E_{z_{0}},\ldots,E_{z_{j}}]}\sum_{m=j}^{q}e^{-\frac{\beta}{2}[E_{z_{r}},\ldots,E_{z_{q}},E_{z_{j}},E_{z_{m}}]}$ =^$\mathrel{\widehat{=}}\int_{0}^{\beta/2}\tau\langle\Lambda_{0}(\tau)\Lambda_{0}\rangle\mathrm{d}\tau$ (104)
$\frac{\delta_{P_{k}}^{(q)}\Lambda_{k}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=1}^{q-1}{\delta_{P_{l}}^{(j)}}{\Lambda_{l}(z_{j})}\sum_{r=0}^{j-1}e^{-\frac{\beta}{2}[E_{z_{0}},\ldots,E_{z_{j}}]}\sum_{m=j-1}^{q-1}e^{-\frac{\beta}{2}[E_{z_{r}},\ldots,E_{z_{q-1}},E_{z_{m}}]}$ =^$\mathrel{\widehat{=}}\int_{0}^{\beta/2}\tau\langle(\Lambda_{k}D_{k}P_{k})(\tau)(\Lambda_{l}D_{l}P_{l})\rangle\mathrm{d}\tau,$ (105)

### v3

@@ F.2 Estimating $\langle H_{1}\rangle$ and $\langle H_{1}(\tau)H_{1}\rangle$
$\frac{\Lambda_{0}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=0}^{q}\Lambda_{0}(z_{j})e^{-\tau[E_{z_{j}},\ldots,E_{z_{q}}]}e^{-(\beta-\tau)[E_{z_{0}},\ldots,E_{z_{j}}]}$ $\mathrel{\widehat{=}}\langle\Lambda_{0}(\tau)\Lambda_{0}\rangle$ (97)
$\frac{\delta_{P_{k}}^{(q)}\Lambda_{k}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=0}^{q}\Lambda_{0}(z_{j})e^{-\tau[E_{z_{j}},\ldots,E_{z_{q-1}}]}e^{-(\beta-\tau)[E_{z_{0}},\ldots,E_{z_{j}}]}$ $\mathrel{\widehat{=}}\langle(\Lambda_{k}D_{k}P_{k})(\tau)\Lambda_{0}\rangle$ (98)
$\frac{\Lambda_{0}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=1}^{q-1}\delta^{(j)}_{P_{l}}\Lambda_{l}(z_{j})e^{-\tau[E_{z_{j}},\ldots,E_{z_{q}}]}e^{-(\beta-\tau)[E_{z_{0}},\ldots,E_{z_{j-1}}]}$ $\mathrel{\widehat{=}}\langle\Lambda_{0}(\tau)(\Lambda_{l}D_{l}P_{l})\rangle$ (99)
$\frac{\delta_{P_{k}}^{(q)}\Lambda_{k}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=1}^{q-1}{\delta_{P_{l}}^{(j)}}{\Lambda_{l}(z_{j})}e^{-\tau[E_{z_{j}},\ldots,{E_{z_{q-1}}}]}e^{-(\beta-\tau)[E_{z_{0}},\ldots,{E_{z_{j-1}}}]}$ $\mathrel{\widehat{=}}\langle(\Lambda_{k}D_{k}P_{k})(\tau)(\Lambda_{l}D_{l}P_{l})\rangle,$ (100)
where each estimator follows as different special cases of the derivation in Sec. VIII.A in Ref. [19]. As with the $\langle D_{0}(\tau)D_{0}\rangle$ estimator, each of these estimators requires $O(q^{2})$ operations, so the total complexity to evaluate $\langle H_{1}(\tau)H_{1}\rangle$ is simply $O(R^{2}q^{2})$.
Although these estimators are more complicated than the pure diagonal counterparts Section E.2, the big picture for ES and FS is the same. Namely, the only non-trivial $\tau$ dependence is contained in the divided differences of the correlator estimator, for which we can apply the divided difference integral relations we have already derived.
F.3 General ES estimator
@@
$\frac{\Lambda_{0}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=0}^{q}\Lambda_{0}(z_{j})e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}},E_{z_{j}}]}$ $\mathrel{\widehat{=}}\int_{0}^{\beta}\langle\Lambda_{0}(\tau)\Lambda_{0}\rangle\mathrm{d}\tau$ (101)
$\frac{\delta_{P_{k}}^{(q)}\Lambda_{k}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=1}^{q-1}{\delta_{P_{l}}^{(j)}}{\Lambda_{l}(z_{j})}e^{-\beta[E_{z_{0}},\ldots,E_{z_{q-1}}]}$ $\mathrel{\widehat{=}}\int_{0}^{\beta}\langle(\Lambda_{k}D_{k}P_{k})(\tau)(\Lambda_{l}D_{l}P_{l})\rangle\mathrm{d}\tau.$ (102)
@@ F.4 General FS estimator
$\frac{\Lambda_{0}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=0}^{q}\Lambda_{0}(z_{j})\sum_{r=0}^{j}e^{-\frac{\beta}{2}[E_{z_{0}},\ldots,E_{z_{j}}]}\sum_{m=j}^{q}e^{-\frac{\beta}{2}[E_{z_{r}},\ldots,E_{z_{q}},E_{z_{j}},E_{z_{m}}]}$ $\mathrel{\widehat{=}}\int_{0}^{\beta/2}\tau\langle\Lambda_{0}(\tau)\Lambda_{0}\rangle\mathrm{d}\tau$ (104)
$\frac{\delta_{P_{k}}^{(q)}\Lambda_{k}(z)}{e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}}\sum_{j=1}^{q-1}{\delta_{P_{l}}^{(j)}}{\Lambda_{l}(z_{j})}\sum_{r=0}^{j-1}e^{-\frac{\beta}{2}[E_{z_{0}},\ldots,E_{z_{j}}]}\sum_{m=j-1}^{q-1}e^{-\frac{\beta}{2}[E_{z_{r}},\ldots,E_{z_{q-1}},E_{z_{m}}]}$ $\mathrel{\widehat{=}}\int_{0}^{\beta/2}\tau\langle(\Lambda_{k}D_{k}P_{k})(\tau)(\Lambda_{l}D_{l}P_{l})\rangle\mathrm{d}\tau,$ (105)